OpenStudy (anonymous):

If f(x) = −3x − 3 and g(x) = 2x + 1, find (−2). −3 3 −1 1

OpenStudy (unklerhaukus):

incomplete question

OpenStudy (anonymous):

no

OpenStudy (unklerhaukus):

(-2)=-2

OpenStudy (anonymous):

When f(-2) or g(-2)..?

OpenStudy (anonymous):

etc.

OpenStudy (unklerhaukus):

\[f(-2)=−3(-2) − 3\] \[g(-2)=2(-2)+ 1\] \[f\circ g(-2)=f(g(-2))=f(2(-2)+ 1)=−3(2(-2)+ 1) − 3\]

OpenStudy (unklerhaukus):

\[g\circ f(-2)=g(f(-2))=g(-3(-2)-3)=2(-3(-2)-3)+1\]

OpenStudy (anonymous):

idk how to do thtt

OpenStudy (unklerhaukus):

well i dont know what your question is are you looking for f(-2) or g(-2)

OpenStudy (anonymous):

if you are looking for f(-2) just put x=-2 in the equation of f(x) f(-2)=-3(-2)-3=6-3= 3 if you are looking for g(-2) just put x=-2 in the equation of g(x) g(-2)=2(-2)+1=-4+1=-3 if you are looking for function composition then look above, unkle has solved those as well.